OpenStudy (anonymous):

f(x)=2x^2-2x-5 f(x+6)=?

OpenStudy (confluxepic):

Try to plug in x+6.

OpenStudy (confluxepic):

f(x+6)=2(x+6)^2-2(x+6)-5

OpenStudy (anonymous):

well i got to this point 2(x^+12x+36)-2x-12-5

OpenStudy (anonymous):

*x^2 sorry

OpenStudy (confluxepic):

f(x+6)=2(x^2+36)-2x+12-5

OpenStudy (confluxepic):

f(x+6)=2x^2+72-2x+7

OpenStudy (confluxepic):

f(x+6)=2x^2-2x+79

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

hmm it says its wrong